one cup of nuts that contain exactly half peanuts and half cashew is added to a bowl of nuts that is exactly 1/3 peanuts,1/3 cashews and 1/3 almonds.This result in a three cup mixture of nuts.What fraction of the new nut mixture is peanuts?
Answer by fcabanski(1391)

If one cup is added to the bowl, and the result is three cups, then the bowl originally held two cups.


1/3 of that 2 cups was peanuts.
1/3 * 2 = 2/3 cup of peanuts in the bowl.


The new mixture added is one cup. Half of that cup is peanuts.
1/2 * 1 = 1/2 cup of peanuts.


The new mixture contains 1/2 + 2/3 = 3/6 + 4/6 = 7/6 cups of peanuts.


It's 7/6 cups of peanuts in 3 cups total = (7/6) / 3 = 7/6 * 1/3 = 7/18 of the new mixture is peanuts.

If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($30/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($3.50-$5.50 per problem). Contact me at fcabanski@hotmail.com